Exercise 5-14 Break-Even and Target Profit Analysis [LO5-3, LO5-4, LO5-5, LO5-6]

Lindon Company is the exclusive distributor for an automotive product that sells for $37.00 per unit and has a CM ratio of 39%. The company’s fixed expenses are $432,900 per year. The company plans to sell 31,000 units this year.


Required:

1.

What are the variable expenses per unit? (Round your answer to 2 decimal places.)

2. Use the equation method:
a.

What is the break-even point in unit sales and in dollar sales? (Do not round intermediate calculations.)

break even point in unit sales?

break-even point in dollar sales?

b.

What amount of unit sales and dollar sales is required to earn an annual profit of $72,150? (Do not round intermediate calculations.)

sales level in units?

sales level in dollars?

c.

Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $4.80 per unit. What is the company’s new break-even point in unit sales and in dollar sales? (Do not round intermediate calculations. Round up break even point answers to the nearest whole number.)

new break even point in unit sales?

new break even point in dollar sales?

3. Repeat (2) above using the formula method.
a.

What is the break-even point in unit sales and in dollar sales? (Do not round intermediate calculations.)

break even point in unit sales?

break even point in dollar sales?

b.

What amount of unit sales and dollar sales is required to earn an annual profit of $72,150? (Do not round intermediate calculations.)

sales level in units?

sales level in dollars?

c.

Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $4.80 per unit. What is the company’s new break-even point in unit sales and in dollar sales? (Do not round intermediate calculations. Round up break even point answers to the nearest whole number.)

new break even point in unit sales?

new break even point in dollar sales?
Answer #1

Selling price -$37 per unit

Contribution margin-39%

Variable cost ratio=1- CM

=61%

2)Selling price $37 per unit

-variable cost 22.57 per unit(61% of 37)

contribution -$14.43 per unit

2 a) break even point in unit sales= fixed cost/ contribution per unit

=$432,900 per year/$14.43 per unit

=30,000 units

break-even point in dollar sales= break even units * selling price price per unit

=30,000*37

=$1110000

2 b) desired profit-$72,150

fixed cost-$432,900

contribution per unit-$14.43 per unit

sales level in units-fixed cost +desired profit/contribution per unit

432900+72150/14.43

=35000 units

sales level in dollars= 35000*37

=$1295000

2 c and 3 a are same) selling price-$37 per unit

variable cost -$17.77 (22.57-4.80)

contribution-$19.23 per unit

new break even point in unit sales- fixed cost/revised contribution per unit

  432,900/19.23

=22512 units

new break even point in dollar sales-22512 units*37

=$832944

3 b) desired profit-$72,150

fixed cost-$432,900

revised contribution -$19.23 per unit

sales level in units-fixed cost +desired profit/ revised contribution per unit

432900+72150/19.23

=26264 units

sales level in dollars= 26264*37

=$971768

  

3 c is same as 2 c